Hi, here's my question: In a sale, a mouse and an elephant together were sold for the price of £97. If the elephant cost £27 more than the mouse. How much did the mouse cost? Thank you for helping me!
Answer by ramkikk66(644)

In a sale, a mouse and an elephant together were sold for the price of £97. If the elephant cost £27 more than the mouse. How much did the mouse cost?
Ans:
Let cost of mouse be M pounds. Then cost of elephant = M + 27
Total cost = M + M + 27 = 97
i.e. 2M = 70 or M = 35

Cost of mouse = 35 pounds, and elephant is 62 pounds.

Hope this helps :)